a rectangular book is 7 cm longer than it is wide. its area is 144 cm find the length and width.
Answer by jim_thompson5910(28715)

A = LW

144 = x(x+7)

x(x+7) = 144

x^2+7x = 144

x^2+7x-144 = 0

(x+16)(x-9) = 0

x+16 = 0 or x-9 = 0

x = -16 or x = 9

Throw out the negative solution to get the only solution of x = 9

So the width is x = 9 cm and the length is x+7 = 9+7 = 16 cm.